OpenStudy (caboose):

How do you find the unilateral Laplace Transform of x(t)=u(t)-u(t-2)

OpenStudy (anonymous):

look it up in my engineering cheetsheet?

OpenStudy (anonymous):

\[X(s)={1 \over s}- {e ^{-2s} \over s}\]

OpenStudy (caboose):

AnwarA, can you explain how you found that? did you just use a laplace transform table?

OpenStudy (anonymous):

you can use the definition of laplace transform

OpenStudy (anonymous):

lemme see if i can make it readable

OpenStudy (anonymous):

\[\int\limits_{0}^{\infty}e^(-s*t)f(t)dt\]

OpenStudy (anonymous):

well you can easily do it using the Laplace definition or this formula for laplace transform of the unit step function: \[L[u(t-a)]={e ^{-as} \over s}\]

OpenStudy (anonymous):

yea, that didn't work, well, you can ignore when the step function is zero, and integrate where t > 0, or t-2 > 0

OpenStudy (caboose):

Ah okay, thank you both.

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

np